# A New Model-Independent Precision Test of General Relativity using LISA Bright Standard Sirens

### v1

@@ 3.3 LISA Source Parameter Estimation
Figure 5: This plot presents the estimated values and uncertainties for the component masses ($m_{1}$ and $m_{2}$), luminosity distance, and inclination angle for LISA sources at a redshift of $z=2.4$. These distributions highlight the precision achievable with LISA’s observational capabilities, underscoring the potential for profound cosmological insights into the properties of binary systems in the early universe. The injected values are as follows: black hole masses $m_{1}=5\times 10^{5}M_{\odot}$, $m_{2}=1\times 10^{5}M_{\odot}$, luminosity distance (denoted by dist) $6.137\times 10^{26}$ meters (19.87Gpc), and inclination angle (denoted by $\iota$) of 1.05 radians.
@@
Figure 6: This figure illustrates the luminosity distance errors as a function of redshift for LISA sources, enhanced to include detailed error analyses. Each data point represents the measured luminosity distance along with the total uncertainty, derived from two specific sources of errors: GW source parameter estimation errors and WL errors. The top subplot shows these absolute errors, highlighting the precision and challenges LISA faces in distance estimation across various redshifts. The lower subplot displays the relative errors, providing insight into how the precision of measurements varies with increasing redshift.
@@
To study the GW events detected by the LISA, it is essential to estimate the parameters of these sources. The process of parameter estimation for LISA sources is quite different from that used for ground-based observatories like LIGO, Virgo, and KAGRA [117]. For our analysis of LISA sources, we employ the BBHx [118] Python package, which is specifically designed for this purpose. Our methodology involves generating waveforms using BBHx [118], which incorporates the IMRPhenomD [20, 119] waveform approximation model with higher order modes. We perform this waveform generation over a one-year observation period. Subsequently, we use BBHx [118] to calculate the likelihood of the generated waveforms. For the estimation of the posterior distributions of key parameters for this study specifically, the masses of the black holes ($m_{1}$ and $m_{2}$), the luminosity distance ($D_{l}$), and the inclination angle ($i$) we constrain the priors of all other parameters to delta functions. Finally, for the parameter estimation, we employ the Dynesty [120] sampler. Among these, the posterior of $D_{l}$ is particularly crucial for our study, as it will be used for the inference of $\mathcal{F}(z)$.
@@
The accuracy of luminosity distance measurements in GW astronomy is influenced not only by the sensitivity and configuration of the detector but also significantly by weak lensing, especially for sources detected at high redshifts by LISA [121]. Gravitational lensing affects GWs similarly to EM radiation. For GW events expected from large redshifts ($z>1$), weak lensing is a common occurrence, along with the occasional strongly-lensed sources. A lens with magnification $\mu$ modifies the observed luminosity distance to $\frac{D_{L}}{\sqrt{\mu}}$, introducing a systematic error $\Delta D_{L}/D_{L}=1-\frac{1}{\sqrt{\mu}}$ [122]. This lensing error, when convolved with the expected magnification distribution $p(\mu)$ from a standard $\Lambda$CDM model, significantly influences the overall measurement accuracy. In this study, we utilize the following model to estimate the error due to weak lensing [123]
$\frac{\sigma_{WL}}{d_{L}}=\frac{0.096}{2}\left(\frac{1-(1+z)^{-0.62}}{0.62}\right)^{2.36}.$ (3.10)
@@
$\sigma_{D_{L}}^{2}=\sigma_{GW}^{2}+\sigma_{WL}^{2},$ (3.11)
where $\sigma_{GW}$ is the uncertainty derived from the detector’s setup during parameter estimation, and $\sigma_{WL}$ accounts for the lensing effects. This comprehensive approach ensures a more accurate understanding of the intrinsic properties of GW sources.
The figure 5 shows the estimated values and uncertainties for component masses, luminosity distance, and inclination angle of LISA sources at a redshift of z=2.4. This highlights the potential of LISA to offer precise insights into the cosmological properties of binary systems from the early universe. The figure 6 presents the luminosity distance errors as a function of redshift for LISA sources, showcasing the variability and precision of measurements across different redshifts. Each point in the plot represents the measured luminosity distance at various redshifts, with total uncertainty, derived from two specific sources of errors: GW source parameter estimation error and WL error. This figure underscores the advanced capabilities of LISA in probing the universe’s most distant reaches, illustrating the impact of different sources of uncertainty on measurement accuracy, and paving the way for profound cosmological discoveries.
@@ 4 Identification of EM Counterparts for LISA Sources
In the near-infrared spectrum, the ELT is anticipated to be a vital tool for redshift measurements of host galaxies. ELT’s MICADO spectrograph will enable the precise measurement of redshifts in the IYJHK bands. The detection threshold for ELT’s spectroscopy is set at an apparent magnitude of 27.2, while for imaging, it extends to 31.3. For galaxies brighter than 27.2, the redshift measurement error is expected to be extremely small, allowing for highly accurate determinations. For fainter galaxies, between magnitudes 27.2 and 31.3, the precision of redshift measurements varies depending on the actual redshift of the source. High-redshift sources, those beyond a redshift of $z=5$, can be identified by the Lyman-alpha break with a redshift error of around 0.2. For sources with redshifts between 0.5 and 5, the Balmer break can be used for redshift identification; however, the absence of observations in the optical and UV parts of the spectrum increases the redshift error to about 0.5 [141, 142, 143, 94].
Collectively, these strategies and instruments provide a comprehensive framework for the detection and study of the EM counterparts of LISA sources. By combining data across optical, radio, X-ray, and near-infrared wavelengths, a multi-faceted understanding of the processes and environments surrounding merging MBBHs and other GW sources can be gained. This multi-wavelength approach is crucial for the advancement of our knowledge of the Universe and for maximizing the revolutionary capabilities of the LISA mission.
@@ 5 Baryon Acoustic Oscillation Scale from galaxy power spectrum
BAO are patterns resembling wrinkles in the distribution density of galaxies across the universe, originating from acoustic density waves in the early universe’s primordial plasma[60, 144]. These waves were propelled by the interplay between radiation pressure and gravitational forces. As the universe expanded and cooled, leading to the decoupling of photons from baryons, a characteristic scale the sound horizon at the drag epoch ($z_{d}$) denoted by $r_{s}$ left a discernible imprint on the galaxy distribution[145, 146, 147, 148, 149].
@@ 5.1 Inference of BAO from galaxy two-point angular correlation function
The imprint manifests as a distinct peak within the two-point angular correlation function (2PACF) of galaxies, $w(\theta)$, which offers a method to measure the universe’s expansion rate across its history. To detect the BAO feature within $w(\theta)$, it is modeled using a combination of a power-law and a Gaussian component, represented by the following equation[150, 151, 152]
$w(\theta,z)=a_{1}+a_{2}\theta^{k}+a_{3}\exp\left(-\frac{(\theta-\theta_{\text{FIT}}(z))^{2}}{\sigma_{\theta}^{2}}\right),$ (5.1)
where $a_{i}$ are parameters representing the fit, $\sigma_{\theta}$ is the width of the BAO peak, and $\theta_{\text{FIT}}$ indicates the angular position of the BAO feature. This method offers a robust standard ruler that enables independent estimates of the angular diameter distance, $D_{a}(z)$, further elucidating the dynamics of the universe’s expansion[153, 154, 155, 156, 157].
Figure 7: This graph displays the relationship of the 2PACF ($\theta^{2}w(\theta)$) as it varies with $\theta$ at a particular redshift, z=1.0, derived from the non linear matter power spectrum. A notable peak appears near $\theta\sim 2.5$, which signifies the detection of the characteristic BAO scale, denoted as $\theta_{BAO}$.
@@ 5.2 Modeling the two-point angular correlation function
The theoretical representation of the 2PACF, $w(\theta)$, is formulated as
$w(\theta)=\sum_{l\geq 0}\left(\frac{2l+1}{4\pi}\right)P_{l}(\cos(\theta))C_{l},$ (5.2)
where $P_{l}$ denotes the Legendre polynomial of the $l^{th}$ order, and $C_{l}$ represents the angular power spectrum. The latter is derived from the three-dimensional matter power spectrum $\mathcal{P}(k)$ through the integral[60, 144]
$C_{l}=\frac{1}{2\pi^{2}}\int 4\pi k^{2}\mathcal{P}(k)\psi_{l}^{2}(k)e^{-k^{2}/k_{eff}^{2}},$ (5.3)
where $\psi_{l}(k)$ is defined as:
@@
Figure 7 showcases the variation of $\theta^{2}w(\theta)$ with angular separation at a fixed redshift $z=1.0$. A marked peak at approximately $\theta\sim 2.5$ reveals the BAO scale, central to our analysis. Furthermore, the covariance of $w(\theta)$, expressed as $Cov_{\theta\theta^{\prime}}$, is computed as:
$Cov_{\theta\theta^{\prime}}=\frac{2}{f_{sky}}\sum_{l\geq 0}\frac{2l+1}{(4\pi)^{2}}P_{l}(\cos(\theta))P_{l}(\cos(\theta^{\prime}))\left(C_{l}+\frac{1}{n}\right)^{2},$ (5.5)
@@
where $1/n$ signifies the shot noise linked to the galaxy’s number density per steradian, and $f_{sky}$ is the fraction of the sky surveyed[162, 163]. A depiction (Figure 8) outlines how the covariance matrix evolves with angular separations at a given redshift(z=1), indicating a growing correlation as $\theta$ and $\theta^{\prime}$ increase. We have included the complete covariance matrix in the analysis, as there are sufficient correlations between different scales.
@@
Figure 8: This graph presents the covariance matrix ($\mathrm{Cov_{\theta\theta^{\prime}}}$), computed from the nonlinear matter power spectrum, plotted against $\theta$ and $\theta^{\prime}$ at a specific redshift, z=1.0. The plot clearly shows that the values of the covariance matrix rise as both $\theta$ and $\theta^{\prime}$ increase, suggesting a greater error at larger angular scales.
@@
Figure 9: This figure illustrates the evolution of the BAO scale and its associated error, primarily due to inaccuracies in photometric redshift (photo-z). For LSST’s planned photo-z error, we show the error on BAO measurements from z=1 to z=3 in cyan. It also depicts the BAO measurement error expected from DESI up to redshift 3.7 (shown in black). Although there are no currently planned instruments to measure BAO beyond redshift 3.7, we extend the BAO measurements up to redshift 6 using the same photo-z error anticipated for Vera Rubin Observatory(in red). This visualization provides essential insights into the variability of the BAO scale and its error with increasing redshift, highlighting a rise in relative error as the standard deviation of the selection function increases with redshift.
@@
Figure 9 elaborates on the BAO scale’s variation and associated error across different redshifts, emphasizing the influence of photo-z errors. For LSST’s planned photo-z error, we illustrate the error on BAO measurements from z=1 to z=3 in cyan, showing how the expected errors will affect the precision of BAO measurements. Vera Rubin Observatory[136, 160], along with projects like Euclid [164], will significantly enhance our understanding of BAO by surveying a total area of 18,000 square degrees, extending precise measurements to higher redshifts. The figure also depicts the BAO measurement error anticipated from Dark Energy Spectroscopic Instrument(DESI) [165] up to redshift 3.7, shown in black, demonstrating the expected performance of DESI in this redshift range. The DESI is a notable five-year terrestrial survey focusing on exploring BAO and the evolution of cosmic structures through redshift-space distortions. Covering 14,000 square degrees of the sky, corresponding to a sky fraction ($f_{\text{sky}}$) of 0.3, DESI aims for an accuracy better than 0.5% within the redshift intervals $3.70.0<z<$ 3.7. Despite the lack of currently planned instruments to measure BAO beyond redshift 3.7, we extend the BAO measurements up to redshift 6 using the same photo-z error anticipated for LSST, shown in red. This extended range provides a hypothetical look at the potential capabilities of future surveys and their ability to probe deeper into the universe. This visualization provides essential insights into the variability of the BAO scale and its error with increasing redshift, highlighting a rise in relative error as the standard deviation of the selection function increases with redshift. With the capability of LISA in probing the high redshift Universe, this plot shows that future galaxy surveys with capabilities of measuring BAO beyond $z=3.5$ can bring interesting tests of fundamental physics explored in this analysis.

### v2

@@ 3.3 LISA Source Parameter Estimation
Figure 6: This plot presents the estimated values and uncertainties for the component masses ($\mathrm{m_{1}}$ and $\mathrm{m_{2}}$), luminosity distance, and inclination angle for LISA source at a redshift of $\mathrm{z=2.4}$. These distributions highlight the precision achievable with LISA’s observational capabilities, underscoring the potential for profound cosmological insights into the properties of binary systems in the early universe. The injected values are as follows: black hole masses $\mathrm{m_{1}=5\times 10^{5}M_{\odot}}$, $\mathrm{m_{2}=1\times 10^{5}M_{\odot}}$, luminosity distance (denoted by dist) $\mathrm{6.137\times 10^{26}}$ meters (19.87 Gpc), and inclination angle (denoted by inc) of 1.05 radians.
@@
Figure 7: This figure illustrates the luminosity distance errors as a function of redshift for LISA sources, enhanced to include detailed error analyses. Each data point represents the measured luminosity distance along with the total uncertainty, derived from two specific sources of errors: GW source parameter estimation errors and WL errors. The error bars correspond to $1$-$\sigma$ uncertainty. The figure highlights the precision and challenges LISA faces in distance estimation across various redshifts and shows how the precision of measurements varies with increasing redshift.
@@
To study the GW events detected by the LISA, it is essential to estimate the parameters of these sources. The process of parameter estimation for LISA sources is quite different from that used for ground-based observatories like LIGO, Virgo, and KAGRA [120]. For our analysis of LISA sources, we employ the BBHx Python package[121] , which is specifically designed for this purpose. Our methodology involves generating waveforms using BBHx [121], which incorporates the IMRPhenomD [20, 122] waveform approximation model. We obtain the waveform for a one-year observation period of the LISA sources. Subsequently, we use BBHx [121] to calculate the likelihood of the generated waveforms. For the estimation of the posterior distributions of key parameters for this study specifically, the masses of the black holes ($\mathrm{m_{1}}$ and $\mathrm{m_{2}}$), the luminosity distance ($\mathrm{D_{L}}$), and the inclination angle ($\mathrm{i}$) we constrain the priors of all other parameters to delta functions. Finally, for the parameter estimation, we employ the Dynesty [123] sampler. Among these, the posterior of $\mathrm{D_{L}}$ is particularly crucial for our study, as it will be used for the inference of $\mathrm{\mathcal{F}(z)}$.
@@
The accuracy of luminosity distance measurements in GW astronomy is influenced not only by the sensitivity and configuration of the detector but also significantly by weak lensing, especially for sources detected at high redshifts by LISA [124]. Gravitational lensing affects GWs similarly to EM radiation. For GW events expected from large redshifts ($\mathrm{z>1}$), weak lensing is a common occurrence, along with the occasional strongly-lensed sources. A lens with magnification $\mu$ modifies the observed luminosity distance to $\mathrm{\frac{D_{L}}{\sqrt{\mu}}}$, introducing a systematic error $\mathrm{\Delta D_{L}/D_{L}=1-\frac{1}{\sqrt{\mu}}}$ [125]. This lensing error, when convolved with the expected magnification distribution $\mathrm{p(\mu)}$ from a standard $\mathrm{\Lambda}$CDM model, significantly influences the overall measurement accuracy. In this study, we utilize the following model to estimate the error due to weak lensing [126]
$\mathrm{\frac{\sigma_{WL}}{D_{L}}=\frac{0.096}{2}\left(\frac{1-(1+z)^{-0.62}}{0.62}\right)^{2.36}}.$ (3.11)
@@
$\mathrm{\sigma_{D_{L}}^{2}=\sigma_{GW}^{2}+\sigma_{WL}^{2}},$ (3.12)
where $\mathrm{\sigma_{GW}}$ is the uncertainty derived from the detector’s setup during parameter estimation, and $\mathrm{\sigma_{WL}}$ accounts for the lensing effects. This comprehensive approach ensures a more accurate understanding of the intrinsic properties of GW sources.
The figure 6 shows the estimated values and uncertainties for component masses, luminosity distance, and inclination angle of LISA source at a redshift of z=2.4. This highlights the potential of LISA to offer precise insights into the cosmological properties of binary systems from the early universe. The figure 7 presents the luminosity distance errors as a function of redshift for LISA sources, showcasing the variability and precision of measurements across different redshifts. Each point in the plot represents the measured luminosity distance at various redshifts, with total uncertainty, derived from two specific sources of errors: GW source parameter estimation error and WL error. This figure underscores the advanced capabilities of LISA in probing the universe’s most distant reaches, illustrating the impact of different sources of uncertainty on measurement accuracy, and paving the way for profound cosmological discoveries.
@@ 4 Identification of EM Counterparts for LISA Sources
In the near-infrared spectrum, the ELT is anticipated to be a vital tool for redshift measurements of host galaxies. ELT’s MICADO spectrograph will enable the precise measurement of redshifts in the IYJHK bands. The detection threshold for ELT’s spectroscopy is set at an apparent magnitude of 27.2, while for imaging, it extends to 31.3. For galaxies brighter than 27.2, the redshift measurement error is expected to be extremely small, allowing for highly accurate determinations. For fainter galaxies, between magnitudes 27.2 and 31.3, the precision of redshift measurements varies depending on the actual redshift of the source. High-redshift sources, those beyond a redshift of $\mathrm{z=5}$, can be identified by the Lyman-alpha break with a redshift error of around 0.2. For sources with redshifts between 0.5 and 5, the Balmer break can be used for redshift identification; however, the absence of observations in the optical and UV parts of the spectrum increases the redshift error to about 0.5 [144, 145, 146, 94]. Collectively, these strategies and instruments provide a comprehensive framework for the detection and study of the EM counterparts of LISA sources. By combining data across optical, radio, X-ray, and near-infrared wavelengths, a multi-faceted understanding of the processes and environments surrounding merging MBBHs and other GW sources can be gained. This multi-wavelength approach is crucial for the advancement of our knowledge of the Universe and for maximizing the revolutionary capabilities of the LISA mission.
@@ 5 Baryon Acoustic Oscillation Scale from galaxy power spectrum
BAO are patterns resembling wrinkles in the distribution density of galaxies across the universe, originating from acoustic density waves in the early universe’s primordial plasma[60, 147]. These waves were propelled by the interplay between radiation pressure and gravitational forces. As the universe expanded and cooled, leading to the decoupling of photons from baryons, a characteristic scale the sound horizon at the drag epoch ($\mathrm{z_{d}}$) denoted by $\mathrm{r_{s}}$ left a discernible imprint on the galaxy distribution[148, 149, 150, 151, 152].
@@ 5.1 Inference of BAO from galaxy two-point angular correlation function
The imprint manifests as a distinct peak within the two-point angular correlation function (2PACF) of galaxies, $\mathrm{w(\theta)}$, which offers a method to measure the universe’s expansion rate across its history. To detect the BAO feature within $\mathrm{w(\theta)}$, it is modeled using a combination of a power-law and a Gaussian component, represented by the following equation[153, 154, 155]
$\mathrm{w(\theta,z)=a_{1}+a_{2}\theta^{k}+a_{3}\exp\left(-\frac{(\theta-\theta_{\text{FIT}}(z))^{2}}{\sigma_{\theta}^{2}}\right)},$ (5.1)
where $a_{i}$ are parameters representing the fit, $\mathrm{\sigma_{\theta}}$ is the width of the BAO peak, and $\mathrm{\theta_{FIT}}$ indicates the angular position of the BAO feature. This method offers a robust standard ruler that enables independent estimates of the angular diameter distance, $\mathrm{D_{a}(z)}$, further elucidating the dynamics of the universe’s expansion[156, 157, 158, 159, 160].
Figure 8: This graph displays the relationship of the 2PACF ($\mathrm{\theta^{2}w(\theta)}$) as it varies with $\mathrm{\theta}$ at a particular redshift, z=1.0, derived from the non linear matter power spectrum. A notable peak appears near $\mathrm{\theta\sim 2.5}$, which signifies the detection of the characteristic BAO scale, denoted as $\mathrm{\theta_{BAO}}$.
@@ 5.2 Modeling the two-point angular correlation function
The theoretical representation of the 2PACF, $\mathrm{w(\theta)}$, is formulated as
$\mathrm{w(\theta)=\sum_{l\geq 0}\left(\frac{2l+1}{4\pi}\right)P_{l}(\cos(\theta))C_{l}},$ (5.2)
where $\mathrm{P_{l}}$ denotes the Legendre polynomial of the $\mathrm{l^{th}}$ order, and $\mathrm{C_{l}}$ represents the angular power spectrum. The latter is derived from the three-dimensional matter power spectrum $\mathrm{\mathcal{P}(k)}$ through the integral[60, 147]
$\mathrm{C_{l}=\frac{1}{2\pi^{2}}\int 4\pi k^{2}\mathcal{P}(k)\psi_{l}^{2}(k)e^{-k^{2}/k_{eff}^{2}}},$ (5.3)
where $\mathrm{\psi_{l}(k)}$ is defined as:
$\mathrm{\psi_{l}(k)=\int dz\phi(z)j_{l}(kr(z))},$ (5.4)
@@
Figure 8 showcases the variation of $\mathrm{\theta^{2}w(\theta)}$ with angular separation at a fixed redshift $\mathrm{z=1.0}$. A marked peak at approximately $\mathrm{\theta\sim 2.5}$ reveals the BAO scale, central to our analysis. Furthermore, the covariance of $\mathrm{w(\theta)}$, expressed as $\mathrm{Cov_{\theta\theta^{\prime}}}$, is computed as:
$\mathrm{Cov_{\theta\theta^{\prime}}=\frac{2}{f_{sky}}\sum_{l\geq 0}\frac{2l+1}{(4\pi)^{2}}P_{l}(\cos(\theta))P_{l}(\cos(\theta^{\prime}))\left(C_{l}+\frac{1}{n}\right)^{2}},$ (5.5)
@@
where $\mathrm{1/n}$ signifies the shot noise linked to the galaxy’s number density per steradian, and $\mathrm{f_{sky}}$ is the fraction of the sky surveyed[165, 166]. A depiction (Figure 9) outlines how the covariance matrix evolves with angular separations at a given redshift(z=1), indicating a growing correlation as $\mathrm{\theta}$ and $\mathrm{\theta^{\prime}}$ increase. We have included the complete covariance matrix in the analysis, as there are sufficient correlations between different scales.
@@
Figure 9: This graph presents the covariance matrix ($\mathrm{Cov_{\theta\theta^{\prime}}}$), computed from the nonlinear matter power spectrum, plotted against $\mathrm{\theta}$ and $\mathrm{\theta^{\prime}}$ at a specific redshift, $\mathrm{z=1.0}$. The plot clearly shows that the values of the covariance matrix rise as both $\mathrm{\theta}$ and $\mathrm{\theta^{\prime}}$ increase, suggesting a greater error at larger angular scales.
@@
Figure 10: This figure illustrates the evolution of the BAO scale and its associated error, primarily due to inaccuracies in photometric redshift (photo-z). For LSST’s planned photo-z error, we show the error on BAO measurements from $\mathrm{z=1}$ to $\mathrm{z=3}$ in cyan. It also depicts the BAO measurement error expected from DESI up to redshift 3.7 (shown in black). Although there are no currently planned instruments to measure BAO beyond redshift 3.7, we extend the BAO measurements up to redshift 6 using the same photo-z error anticipated for Vera Rubin Observatory(in red). This visualization provides essential insights into the variability of the BAO scale and its error with increasing redshift, highlighting a rise in relative error as the standard deviation of the selection function increases with redshift.
@@
Figure 10 elaborates on the BAO scale’s variation and associated error across different redshifts, emphasizing the influence of photo-z errors. For LSST’s planned photo-z error, we illustrate the error on BAO measurements from $\mathrm{z=1}$ to $\mathrm{z=3}$ in cyan, showing how the expected errors will affect the precision of BAO measurements. Vera Rubin Observatory[139, 163], along with projects like Euclid [167], will significantly enhance our understanding of BAO by surveying a total area of 18,000 square degrees, extending precise measurements to higher redshifts. The figure also depicts the BAO measurement error anticipated from Dark Energy Spectroscopic Instrument(DESI) [168] up to redshift 3.7, shown in black, demonstrating the expected performance of DESI in this redshift range. The DESI is a notable five-year terrestrial survey focusing on exploring BAO and the evolution of cosmic structures through redshift-space distortions. Covering 14,000 square degrees of the sky, corresponding to a sky fraction ($\mathrm{f_{\text{sky}}}$) of 0.3, DESI aims for an accuracy better than 0.5% within the redshift intervals $\mathrm{0.0<z<3.7}$. Despite the lack of currently planned instruments to measure BAO beyond redshift 3.7, we extend the BAO measurements up to redshift 6 using the same photo-z error anticipated for LSST, shown in red. This extended range provides a hypothetical look at the potential capabilities of future surveys and their ability to probe deeper into the universe. This visualization provides essential insights into the variability of the BAO scale and its error with increasing redshift, highlighting a rise in relative error as the standard deviation of the selection function increases with redshift. With the capability of LISA in probing the high redshift Universe, this plot shows that future galaxy surveys with capabilities of measuring BAO beyond $\mathrm{z=3.5}$ can bring interesting tests of fundamental physics explored in this analysis.